# Mechanically Interlocked Polymers in Dilute Solution under Shear and Extensional Flows: A Brownian Dynamics Study

### v2

@@ III.1 Polyrotaxanes: Extensional Flow
At the lower end of Wi range, the behavior of polyrotaxanes diverges significantly from linear polymers. Fig. 2a shows that adding rings to linear polymers increases the fractional extension in the flow direction. This low Wi effect is largely equilibrium in nature, as demonstrated in SI Tables A.1 and A.2, which give the equilibrium radius of gyration $R_{g}$, asphericity Theodorou and Suter (1985), and prolateness Vatin et al. (2024) of linear polymers and polyrotaxanes. Shorter polyrotaxanes with a high density of mechanical bonds are more extended (higher $R_{g}$), less spherical (higher asphericity), and more rod-like (higher prolateness) than linear polymers with a similar backbone length. This is in agreement with experimental works on PEG-based polyrotaxanes where more rings elongate and rigidify the backbone Qiu et al. (2020). In general, both rings and backbones are rigidified after threading as rotaxanes, even when the starting components are flexible because the intermolecular and steric interactions reduce the conformational space and increase barriers Fadler and Flood (2022). The differences in extension become less pronounced for longer backbones or fewer mechanical bonds, which is also reflected in Fig. 2a. The low Wi differences, then, can be understood as a difference in equilibrium shape that will align with weak flows. At higher Wi values, differences due to interlocked rings become negligible as both linear and polyrotaxane polymers reach fully extended backbone configurations. Examples of linear polymers and polyrotaxanes in these steady-state stretched configurations are shown in SI Fig. C.1.
The first normal stress difference $(N_{1})$ is lower in polyrotaxanes than linear polymers across the range of simulated Wi values, as shown in Fig. 2b. Polymers with more rings exhibit lower $N_{1}$ values compared with linear polymers, demonstrating that mechanical bonds lead to weaker elastic behavior. This weaker elastic behavior is notably similar to bottlebrush polymers Dutta and Sing (2024), with increasing bristle length playing a similar role to backbone ring density in polyrotaxanes.
@@ III.2 Polyrotaxanes: Shear Flow
Polymers in shear flow undergo continuous stretch-tumble cycles De Gennes (1974); Smith et al. (1999); Teixeira et al. (2005) and the fractional extension values reported in this work are averaged over those cycles. Example configurations of a polyrotaxane (and other MIP types) during these cycles are shown in SI Fig. C.1. According to Fig. 3a, fractional extension in flow direction ($\langle X_{bb}\rangle/L$) in polyrotaxanes is higher than the linear polymers with the same number of backbone beads across the Wi range tested. A greater number of rings result in higher extension and more deviation from linear polymers. Similar to low Wi extensional flow, the ring-backbone interactions increase rigidity and cause the molecules to exhibit more elongated configurations on average, resulting in comparatively higher extension in flow direction. In real rotaxanes, this increased rigidity can be caused by hydrogen bonding between the ring and backbone, $\pi$-$\pi$ stacking between rings, metal coordination, steric interactions between rings, or electrostatic repulsions Fadler and Flood (2022).
@@
Polyrotaxanes exhibit higher shear viscosity $(\eta)$ than their linear counterparts, with larger number of threaded rings leading to further rise in viscosity, as shown in Fig. 3b. The size of a polymer in solution has a direct relationship with the viscosity of the solution Krigbaum and Flory (1953); Hiemenz and Lodge (2007); Rubinstein and Colby (2003). As evident from SI Tables A.1 and A.2, the radius of gyration of polyrotaxanes is higher than linear polymers with the same backbone length, with additional threaded rings leading to larger polymer size. For example, a 120 bead linear polymer has a radius of gyration of 15.86, compared to a polyrotaxane with 24 rings on a 120 bead backbone, which has a radius of gyration of 23.80.
@@
Similar to linear polymers, polyrotaxanes show shear thinning, although to a lesser degree, as evident from the power-law exponent $\beta$ values shown in SI Table E.1. Additionally, polyrotaxanes with more rings have lower $\beta$ values. For example, a linear polymer with 80 beads has $\beta=-0.380$ (in agreement with the previously reported value for this model Dutta and Sing (2024)) whereas threading 4 or 16 rings on it decreases the $\beta$ value to 0.290-0.290 and 0.225-0.225, respectively, showing that more rings further suppress shear thinning. The weaker shear thinning is due to the polyrotaxanes’ decreased alignment with the flow, demonstrated by the average orientation angle in SI Fig. B.1c. For $\mathrm{Wi}>1$, increasing the number of rings results in less alignment with the flow on average (an angle $\theta$ with greater deviation from $0^{\circ}$) because the rings cannot fully align with the flow due to repulsive interactions within the rings and between the rings and the backbone. This misalignment leads to less relaxation and weaker shear thinning Teixeira et al. (2005); Ryder and Yeomans (2006).
@@
In weak flow ($\mathrm{Wi}\leq 1$), shear forces are not sufficient to induce significant tumbling in either polyrotaxanes or linear polymers, as shown in Fig. 3c. Consequently, both polymers have flat tumbling frequency profiles. As the flow strength increases ($1<\mathrm{Wi}<10$), polyrotaxanes start to show enhanced tumbling compared to the linear counterparts. At $\mathrm{Wi}\geq 10$, the polyrotaxanes show significantly higher tumbling frequency compared with linear polymers. At $\mathrm{Wi}=800$, polyrotaxanes with 40 and 80 backbone beads (with 8 and 16 threaded rings, respectively) exhibit tumbling frequencies 38% and 35% higher than their linear polymer counterparts. Furthermore, within polyrotaxanes, threading more rings on a backbone increases the tumbling frequency. For instance, at $\mathrm{Wi}=10^{3}$, a polyrotaxane with 80 beads in the backbone and 16 threaded rings tumbles 14% more frequently than one with only 4 rings. The enhanced tumbling in polyrotaxanes can be explained by the threaded rings increasing the polymer’s hydrodynamic cross-section along the gradient direction, as shown is SI Fig. B.1a. The expanded profile exposes a larger surface area to shear forces of the solvent, leading to more frequent tumbling events. The polyrotaxanes also have a greater extent in the cross-flow direction, demonstrated in SI Fig. B.1b.
@@
Fig. 3d shows the first normal stress coefficient ($\psi_{1}$) as a function of Wi. At a comparable backbone length, polyrotaxanes have higher $\psi_{1}$ than linear polymers, with more rings further increasing the value. As shown in SI Table E.2, the power-law exponent ($\beta$) in polyrotaxanes is only slightly higher than linear polymers. For example, for a polyrotaxane with 120 beads in the backbone and 24 threaded rings, $\beta=-1.24$ while for a linear polymer with $N=120$, we report $\beta=-1.32$. The latter value is in close agreement with the previous reported value of -1.30 for this model Dutta and Sing (2024).
@@
Figure 4: Daisy chain properties in extensional flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X_{bb}\rangle/L)$, (b) first normal stress difference $(N_{1})$, (c) extensional viscosity $(\eta_{e})$. Linear polymers are shown in shades of gray with darker shades indicating more beads. Daisy chains with 8, 16, and 32 beads per segment (BPS) are shown in blue, green, and red, respectively; darker shades indicate more segments ($N_{\mathrm{seg}}$). The rings in all daisy chains are composed of 8 beads.
@@
Figure 5: Daisy chain properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X_{bb}\rangle/L)$, (b) shear viscosity ($\eta$), (c) tumbling frequency ($\omega\tau_{R}$), and (d) first normal stress coefficient ($\psi_{1}$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Daisy chains with 8, 16, and 32 beads per segment (BPS) are shown in blue, green, and red, respectively; darker shades indicate higher number of segments ($N_{\mathrm{seg}}$). The rings in all daisy chains are composed of 8 beads. Fitted power law exponents for $\mathrm{Wi}>10^{2}$ for $\eta$ and $\psi_{1}$ are reported in SI Tables E.1 and E.2, respectively.
@@ III.3 Daisy Chains: Extensional Flow
From Fig. 4a, at low Wi in extensional flow, daisy chains are more extended than linear polymers at similar $M_{w}$ with higher density of mechanical bonds leading to greater fractional extension. As an example, at Wi=0.1, a daisy chain with 8 beads per segment and 4 segments ($M_{w}=74$) is 72% more extended than a comparable linear polymer ($M_{w}=80$) whereas a daisy chain with 32 beads per segment and 2 segments ($M_{w}=78$), which has lower density of mechanical bonds, is only 22% more extended. This is reflected in the shape parameters of these two polymers in equilibrium, where the chain with more short segments is more prolate (0.74 vs. 0.49) and has higher asphericity (0.70 vs. 0.61), demonstrated in SI Tables A.1 and A.3.
This behavior is similar to backbone rigidifying effect of the rings as seen in the polyrotaxanes. In chains with longer segments, this effect diminishes and the chains become more coiled like a linear polymer. Interestingly, shorter linear segments lead to greater extension at low Wi, but the trend reverses at high Wi where daisy chains with longer linear segments are more extended. In strong flow regime where the chains are far from equilibrium and more extended, the repulsions between rings becomes important and a stronger flow should be needed to stretch daisy chains with more segments. Making segments longer diminishes this effect and the chains become more like a linear polymer.
@@ III.4 Daisy Chains: Shear Flow
As shown in Fig. 5a, extension in the flow direction ($\langle X_{bb}\rangle/L$) has different regimes depending on flow strength. Previous studies on semi-flexible linear polymers show a similar shear response where polymers stretch at low shear rates and collapse at high shear rates Lyulin et al. (1999). At $\mathrm{Wi}<0.8$, $\langle X_{bb}\rangle/L$ remains nearly constant in linear polymers and daisy chains, with daisy chains exhibiting greater extension than linear polymers of similar molecular weight. For example, at $\mathrm{Wi}=10^{-1}$, a daisy chain of 6 linear segments with 8 beads each ($M_{w}=118$) has a fractional extension of 0.16 compared to a linear polymer of $M_{w}=120$ with an extension of 0.09. At $10^{0}<\mathrm{Wi}<10^{2}$, $\langle X_{bb}\rangle/L$ increases and then decreases around $\mathrm{Wi}\approx 10^{2}$. While the conformational change is seen in both linear and daisy chains, the variation is less significant in daisy chains with more mechanical bonds. For example, a daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) has a fractional extension between 0.16 and 0.27 across the Wi range compared to a linear polymer of $M_{w}=120$, which has extensions between 0.08 and 0.029 across the same Wi range. Decreasing the number of segments or increasing the length of the linear segments increases the variation in extension. Also, in daisy chains with the same number of segments, longer segments result in behavior similar to a linear polymer. A terminal regime with increasing extension is seen at very high Wi in linear polymers. However, this behavior could not be confirmed in daisy chains because with our current model, the chains disentangle under very strong flow conditions due to bond crossing. We plan to address this shortcoming in future work.
@@
Fig. 5b shows that the shear thinning degree is lower in daisy chains compared to linear polymers and the power-law exponent $\beta$ (SI Table E.1) is lower in daisy chains with higher mechanical bond density. For example, a daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) has $\beta=-0.225$ while the corresponding linear polymer ($M_{w}=120$) has $\beta=-0.390$, which shows the less pronounced shear-thinning character of the daisy chains. Shear-thinning behavior becomes more linear-like as the daisy chains’ linear segment length increases. Similar to polyrotaxanes, this decreased shear-thinning in daisy chains is caused by their misalignment with the flow direction. SI Fig. B.2c shows that the orientation angle deviates from the flow direction more for daisy chains with shorter linear segments or fewer segments. At a comparable molecular weight, daisy chains also have lower zero-shear viscosity due to their smaller equilibrium size, as shown in SI Table A.3, highlighting another significant deviation from linear polymers. For example, at $\mathrm{Wi}=20$, a daisy chain with 32 beads per segment and 2 segments ($M_{w}=78$) has 42% lower viscosity than the corresponding linear polymer ($M_{w}=80$). It is worth noting that because of the difference in shear-thinning degree between linear and daisy chain polymers, the viscosity difference is more pronounced in weaker flows and diminishes in stronger flows. As an example, a daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) has 520% lower viscosity than the comparable linear polymer ($M_{w}=120$) at $\mathrm{Wi}=0.1$ but the difference decreases at $\mathrm{Wi}=600$ to a less significant value of  27%.
@@
Based on Fig. 5c, at low Wi, the flow is not strong enough to cause significant tumbling in either the daisy chains or linear polymers. At higher Wi (>7>7), daisy chains show significantly more frequent tumbling in shear flow. At $\mathrm{Wi}=200$, a daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) shows 87% higher $\omega\tau_{R}$ than a linear polymer with 120 beads. Similarly, a daisy chain with 32 beads per segment and 2 segments ($M_{w}=78$) has 60% higher $\omega\tau_{R}$ than the linear counterpart with 80 beads. In shear flow, the greater extent of polymers in the gradient direction (shown in SI Fig. B.2a) causes larger shear forces from the flow to be applied to the chains, resulting in more frequent tumbles. The rings in the daisy chains increase the gradient-flow extent. This effect is less present in configurations with lower density of mechanical bonds, making them more linear-like in terms of aspect ratio, showing less tumbling. On the other hand, the role of the rings in the aspect ratio is more pronounced in daisy chains with higher mechanical bond density, causing the chains to experience larger shear forces from the flow, resulting in higher tumbling frequency.
@@
According to Fig. 5d, daisy chains show lower $\psi_{1}$ compared with linear polymers at the same molecular weight, with higher density of mechanical bonds further decreasing $\psi_{1}$ value. As shown in SI Table E.2, the power-law exponent ($\beta$) for daisy chains can be lower or higher than linear polymers with comparable molecular weight depending on the density of mechanical bonds. For example, for a daisy chain with 32 beads per segment and 2 segments ($M_{w}=78$), $\beta=-1.349$ which is lower than the value of -1.239 for a linear polymer with 80 beads. On the other hand, a daisy chain with 8 beads per segment and 4 segments ($M_{w}=74$) has $\beta=-1.105$ due to higher density of mechanical bonds.
@@
Figure 6: Polycatenane properties in extensional flow as a function of Weissenberg number (Wi): (a, b) fractional extension in flow direction $(\langle X_{bb}\rangle/L)$, (c, d) first normal stress difference $(N_{1})$, (e, f) extensional viscosity $(\eta_{e})$. Linear polymers are shown in shades of gray with darker shades indicating more beads. Colors from cold to hot indicate higher number of rings ($N_{\mathrm{rings}}$) with fixed beads per ring ($\mathrm{BPR}=8$). Darker shades of green indicate increasing beads per ring (BPR) in polycatenanes with fixed number of rings ($N_{\mathrm{rings}}=10)$.
@@
Figure 7: Poylcatenane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X_{bb}\rangle/L)$, (b) shear viscosity ($\eta$), (c) tumbling frequency ($\omega\tau_{R}$), and (d) first normal stress coefficient ($\psi_{1}$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Colors from cold to hot indicate higher number of rings ($N_{\mathrm{rings}}$) with fixed beads per ring ($\mathrm{BPR}=8$). Fitted power law exponents for $\mathrm{Wi}>10^{2}$ for $\eta$ and $\psi_{1}$ are reported in SI Tables E.1 and E.2, respectively.
@@
Figure 8: Poylcatenane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X_{bb}\rangle/L)$, (b) shear viscosity ($\eta$), (c) tumbling frequency ($\omega\tau_{R}$), and (d) first normal stress coefficient ($\psi_{1}$). Darker shades of green indicate increasing beads per ring (BPR) in polycatenanes with fixed number of rings ($N_{\mathrm{rings}}=10)$. Fitted power law exponents for $\mathrm{Wi}>10^{2}$ for $\eta$ and $\psi_{1}$ are reported in SI Tables E.1 and E.2, respectively.

### v3

@@ III.1 Polyrotaxanes: Extensional Flow
At the lower end of Wi range, the behavior of polyrotaxanes diverges significantly from linear polymers. Fig. 2a shows that adding rings to linear polymers increases the fractional extension in the flow direction. This low Wi effect is largely equilibrium in nature, as demonstrated in SI Tables A.1 and A.2, which give the equilibrium radius of gyration $R_{g}$, asphericity theodorou1985shape, and prolateness vatin2024conformation of linear polymers and polyrotaxanes. Shorter polyrotaxanes with a high density of mechanical bonds are more extended (higher $R_{g}$), less spherical (higher asphericity), and more rod-like (higher prolateness) than linear polymers with a similar backbone length. This is in agreement with experimental works on PEG-based polyrotaxanes where more rings elongate and rigidify the backbone qiu2020precise. In general, both rings and backbones are rigidified after threading as rotaxanes, even when the starting components are flexible because the intermolecular and steric interactions reduce the conformational space and increase barriers fadler2022rigidity. This leads to a much less pronounced transition between a coiled state and stretched state in polycatenanes, where the stiffening effect of the rings is similar to the reduced coil-stretch transition seen in linear polymers as the backbone becomes stiffer andrews1998configuration. The differences in extension become less pronounced for longer backbones or fewer mechanical bonds, which is also reflected in Fig. 2a. The low Wi differences, then, can be understood as a difference in equilibrium shape that will align with weak flows. At higher Wi values, differences due to interlocked rings become negligible as both linear and polyrotaxane polymers reach more extended backbone configurations. Examples of linear polymers and polyrotaxanes in these steady-state stretched configurations are shown in SI Fig. C.1. We omit extensional flow data for $\mathrm{Wi}>100$ because mechanical bonds in polyrotaxanes fail in this regime as hydrodynamic forces become strong enough to de-thread rings past the end caps.
The first normal stress difference $(N_{1})$ is lower in polyrotaxanes than linear polymers across the range of simulated Wi values, as shown in Fig. 2b. Polymers with more rings exhibit lower $N_{1}$ values compared with linear polymers, demonstrating that mechanical bonds lead to weaker elastic behavior. This weaker elastic behavior is notably similar to bottlebrush polymers dutta2024brownian, with increasing bristle length playing a similar role to backbone ring density in polyrotaxanes.
@@ III.2 Polyrotaxanes: Shear Flow
Polymers in shear flow undergo continuous stretch-tumble cycles de1974coil; smith1999single; teixeira2005shear and the fractional extension values reported in this work are averaged over those cycles. Example configurations of a polyrotaxane (and other MIP types) during these cycles are shown in SI Fig. C.1. According to Fig. 3a, fractional extension in flow direction ($\langle X\rangle/L$) in polyrotaxanes is higher than the linear polymers with the same number of backbone beads across the Wi range tested. A greater number of rings result in higher extension and more deviation from linear polymers. Similar to low Wi extensional flow, the ring-backbone interactions increase rigidity and cause the molecules to exhibit more elongated configurations on average, resulting in comparatively higher extension in flow direction. In real rotaxanes, this increased rigidity can be caused by hydrogen bonding between the ring and backbone, $\pi$-$\pi$ stacking between rings, metal coordination, steric interactions between rings, or electrostatic repulsions fadler2022rigidity.
@@
Polyrotaxanes exhibit higher shear viscosity $(\eta)$ than their linear counterparts, with larger number of threaded rings leading to further rise in viscosity, as shown in Fig. 3b. The size of a polymer in solution has a direct relationship with the viscosity of the solution krigbaum1953molecular; hiemenz2007polymer; rubinstein2003polymer. As evident from SI Tables A.1 and A.2, the radius of gyration of polyrotaxanes is higher than linear polymers with the same backbone length, with additional threaded rings leading to larger polymer size. For example, a 120 bead linear polymer has a radius of gyration of 15.9, compared to a polyrotaxane with 24 rings on a 120 bead backbone, which has a radius of gyration of 23.8.
@@
Similar to linear polymers, polyrotaxanes show shear thinning, although to a lesser degree, as evident from the power-law scaling exponent $\beta$ values shown in SI Table E.1. Additionally, polyrotaxanes with more rings have lower $\beta$ values. For example, a linear polymer with 80 beads has $\beta=-0.380$ in agreement with the previously reported value for this model dutta2024brownian and similar to other dilute linear polymers with excluded volume and hydrodynamic interactions ($\beta=-0.28$moghani2017computationally. Threading 4 or 16 rings onto the backbone decreases the power law scaling value to 0.290-0.290 and 0.225-0.225, respectively, showing that more rings further suppress shear thinning. The weaker shear thinning is due to the polyrotaxanes’ decreased alignment with the flow, demonstrated by the average orientation angle in SI Fig. B.1c. Increasing the number of rings results in less alignment with the flow on average (an angle $\theta$ with greater deviation from $0^{\circ}$) because the rings cannot fully align with the flow due to repulsive interactions within the rings and between the rings and the backbone. This misalignment leads to less relaxation and weaker shear thinning teixeira2005shear; ryder2006shear.
@@
In weak flow ($\mathrm{Wi}\leq 1$), shear forces are not sufficient to induce significant tumbling in either polyrotaxanes or linear polymers, as shown in Fig. 3c. Consequently, both polymers have flat tumbling frequency profiles. As the flow strength increases ($1<\mathrm{Wi}<10$), polyrotaxanes start to show enhanced tumbling compared to the linear counterparts. At $\mathrm{Wi}\geq 10$, the polyrotaxanes show significantly higher tumbling frequency compared with linear polymers. At $\mathrm{Wi}=800$, polyrotaxanes with 40 and 80 backbone beads (with 8 and 16 threaded rings, respectively) exhibit tumbling frequencies 38% and 35% higher than their linear polymer counterparts. Furthermore, increasing the number of threaded rings increases the tumbling frequency. For instance, at $\mathrm{Wi}=10^{3}$, a polyrotaxane with 80 beads in the backbone and 16 threaded rings tumbles 14% more frequently than one with only 4 rings. Threading rings along the backbone also enhances the scaling of the tumbling frequency with Wi. As shown in SI Table E.3, adding 8 rings to a linear polymer with $N=40$ increases the scaling exponent from 0.741 to 0.826. The exponent obtained for linear polymers is consistent with the previously reported 3/4 power-law scaling of tumbling time with shear rate saha2012tumbling. The enhanced tumbling in polyrotaxanes can be explained by the threaded rings increasing the polymer’s hydrodynamic cross-section along the gradient direction, as shown is SI Fig. B.1a. The expanded profile exposes a larger surface area to shear forces of the solvent, leading to more frequent tumbling events. The polyrotaxanes also have a greater extent in the vorticity direction, demonstrated in SI Fig. B.1b.
@@
Fig. 3d shows the first normal stress coefficient ($\psi_{1}$) as a function of Wi. At a comparable backbone length, polyrotaxanes have higher $\psi_{1}$ than linear polymers, with more rings further increasing the value. As shown in SI Table E.2, the power-law exponent ($\beta$) in polyrotaxanes is only slightly higher than linear polymers. For example, for a polyrotaxane with 120 beads in the backbone and 24 threaded rings, $\beta=-1.24$ while for a linear polymer with $N=120$, we report $\beta=-1.32$. The latter value is in close agreement with the previous reported value of -1.30 for this model dutta2024brownian and other excluded volume linear polymer models with hydrodynamic interactions (-1.13) moghani2017computationally. Generally, the power law scaling of $\psi_{1}$ is reduced in polyrotaxanes relative to that of linear polymers with the same backbone size. A similar effect was seen in bottlebrush polymers where side chains are grafted onto a linear backbone dutta2024brownian.
@@
Figure 4: Daisy chain properties in extensional flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X\rangle/L)$, (b) first normal stress difference $(N_{1})$, (c) extensional viscosity $(\eta_{e})$. Linear polymers are shown in shades of gray with darker shades indicating more beads. Daisy chains with 8, 16, and 32 beads per segment (BPS) are shown in blue, green, and red, respectively; darker shades indicate more segments ($N_{\mathrm{seg}}$). The rings in all daisy chains are composed of 8 beads.
@@
Figure 5: Daisy chain properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X\rangle/L)$, (b) shear viscosity ($\eta$), (c) tumbling frequency ($\omega\tau_{R}$), and (d) first normal stress coefficient ($\psi_{1}$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Daisy chains with 8, 16, and 32 beads per segment (BPS) are shown in blue, green, and red, respectively; darker shades indicate higher number of segments ($N_{\mathrm{seg}}$). The rings in all daisy chains are composed of 8 beads. Fitted power law exponents for $\mathrm{Wi}>10^{2}$ for $\eta$ and $\psi_{1}$ are reported in SI Tables E.1 and E.2, respectively.
@@ III.3 Daisy Chains: Extensional Flow
From Fig. 4a, at low Wi in extensional flow, daisy chains are more extended than linear polymers at similar $M_{w}$ with higher density of mechanical bonds leading to greater fractional extension. As an example, at $\mathrm{Wi}=0.1$, a daisy chain with 8 beads per segment and 4 segments ($M_{w}=74$) is 72% more extended than a comparable linear polymer ($M_{w}=80$) whereas a daisy chain with 32 beads per segment and 2 segments ($M_{w}=78$), which has lower density of mechanical bonds, is only 22% more extended. This is reflected in the shape parameters of these two polymers in equilibrium, where the chain with more short segments is more prolate (0.74 vs. 0.49) and has higher asphericity (0.70 vs. 0.61), demonstrated in SI Tables A.1 and A.3.
This behavior is similar to the backbone rigidifying effect of the rings as seen in the polyrotaxanes. In chains with longer segments, this effect diminishes and the chains become more coiled like a linear polymer. Interestingly, shorter linear segments lead to greater extension at low Wi, but the trend reverses at high Wi where daisy chains with longer linear segments are more extended. In the strong flow regime where the chains are far from equilibrium and more extended, the repulsions between rings become important and a stronger flow is be needed to stretch daisy chains with more segments. Making segments longer diminishes this effect and the chains become more like a linear polymer. Overall, a higher density of mechanical bonds “stiffens” the polymer, leading to lower variability in the stretch and a weaker coil-stretch transition. We report extensional flow data up to $\mathrm{Wi}=20$ where the hydrodynamic forces in cause bond crossing between the interlocked rings.
@@ III.4 Daisy Chains: Shear Flow
As shown in Fig. 5a, fraction extension in the flow direction has different regimes depending on flow strength. Previous studies on semi-flexible linear polymers show a similar shear response where polymers stretch at low shear rates and collapse at high shear rates lyulin1999brownian. At $\mathrm{Wi}<0.8$, extension remains nearly constant in linear polymers and daisy chains, with daisy chains exhibiting greater extension than linear polymers of similar molecular weight. For example, at $\mathrm{Wi}=10^{-1}$, a daisy chain of 6 linear segments with 8 beads each ($M_{w}=118$) has a fractional extension of 0.16 compared to a linear polymer of $M_{w}=120$ with an extension of 0.09. At $10^{0}<\mathrm{Wi}<10^{2}$, extension increases and then decreases around $\mathrm{Wi}\approx 10^{2}$. While the conformational change is seen in both linear and daisy chains, the variation is less significant in daisy chains with more mechanical bonds. For example, a daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) has a fractional extension between 0.16 and 0.27 across the Wi range compared to a linear polymer of $M_{w}=120$, which has extensions between 0.08 and 0.029 across the same Wi range. Decreasing the number of segments or increasing the length of the linear segments increases the variation in extension. Also, in daisy chains with the same number of segments, longer segments result in behavior similar to a linear polymer. A terminal regime with increasing extension is seen at very high Wi in linear polymers. However, this behavior could not be confirmed in daisy chains because with our current model the chains disentangle under very strong flow conditions due to bond crossing. We plan to address this shortcoming in future work.
@@
Fig. 5b shows that the shear thinning degree is lower in daisy chains compared to linear polymers and the shear viscosity power-law scaling exponent $\beta$ (SI Table E.1) is lower in daisy chains with higher mechanical bond density. For example, a daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) has $\beta=-0.225$ while the corresponding linear polymer ($M_{w}=120$) has $\beta=-0.390$, which shows the less pronounced shear-thinning character of the daisy chains. Shear-thinning behavior becomes more linear-like as the daisy chains’ linear segment length increases. Similar to polyrotaxanes, this decreased shear-thinning in daisy chains is caused by their misalignment with the flow direction. SI Fig. B.2c shows that the orientation angle deviates from the flow direction more for daisy chains with shorter linear segments or fewer segments. At a comparable molecular weight, daisy chains also have lower zero-shear viscosity due to their smaller equilibrium size, as shown in SI Table A.3, highlighting another significant deviation from linear polymers. For example, at $\mathrm{Wi}=20$, a daisy chain with 32 beads per segment and 2 segments ($M_{w}=78$) has 42% lower viscosity than the corresponding linear polymer ($M_{w}=80$). It is worth noting that because of the difference in shear-thinning degree between linear and daisy chain polymers, the viscosity difference is more pronounced in weaker flows and diminishes in stronger flows. As an example, a daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) has a factor of 5.2 lower viscosity than the comparable linear polymer ($M_{w}=120$) at $\mathrm{Wi}=0.1$ but the difference decreases at $\mathrm{Wi}=600$ to a less significant value of  27%.
@@
Based on Fig. 5c, at low Wi, the flow is not strong enough to cause significant tumbling in either the daisy chains or linear polymers. At higher Wi (>7>7), daisy chains show significantly more frequent tumbling in shear flow. At $\mathrm{Wi}=200$, a daisy chain with 8 beads per segment and 6 segments ($M_{w}=118$) shows 87% higher $\omega\tau_{R}$ than a linear polymer with 120 beads. Similarly, a daisy chain with 32 beads per segment and 2 segments ($M_{w}=78$) has 60% higher $\omega\tau_{R}$ than the linear counterpart with 80 beads. Tumbling frequency also scales more strongly with Wi in daisy chains than linear polymers. As shown in SI table E.3, the scaling exponent of a daisy chain with $M_{w}=118$ is 0.838 compared to 0.732 of the equivalent linear polymer. In fact, a daisy chain of two 16-bead linear segments exhibits a scaling exponent of 0.875, compared to the expected 0.75 exponent of a linear polymer saha2012tumbling. The rings in the daisy chains increase the gradient-direction extent (shown in SI Fig. B.2a) causes larger shear forces from the flow to be applied to the chains, resulting in more frequent tumbles. This effect is less present in configurations with lower density of mechanical bonds, making them more linear-like in terms of aspect ratio, showing less tumbling. On the other hand, the role of the rings in the aspect ratio is more pronounced in daisy chains with higher mechanical bond density, causing the chains to experience larger shear forces from the flow, resulting in higher tumbling frequency.
@@
According to Fig. 5d, daisy chains show lower $\psi_{1}$ compared with linear polymers at the same molecular weight, with higher density of mechanical bonds further decreasing $\psi_{1}$ value. As shown in SI Table E.2, the power-law scaling exponent ($\beta$) for daisy chains can be lower or higher than linear polymers with comparable molecular weight depending on the density of mechanical bonds. For example, a daisy chain with 32 beads per segment and 2 segments ($M_{w}=78$), $\beta=-1.349$ which is lower than the value of -1.239 for a linear polymer with 80 beads. On the other hand, a daisy chain with 8 beads per segment and 4 segments ($M_{w}=74$) has $\beta=-1.105$ due to higher density of mechanical bonds.
@@
Figure 6: Polycatenane properties in extensional flow as a function of Weissenberg number (Wi): (a, b) fractional extension in flow direction $(\langle X\rangle/L)$, (c, d) first normal stress difference $(N_{1})$, (e, f) extensional viscosity $(\eta_{e})$. Linear polymers are shown in shades of gray with darker shades indicating more beads. Colors from cold to hot indicate higher number of rings ($N_{\mathrm{rings}}$) with fixed beads per ring ($\mathrm{BPR}=8$). Darker shades of green indicate increasing beads per ring (BPR) in polycatenanes with fixed number of rings ($N_{\mathrm{rings}}=10)$.
@@
Figure 7: Poylcatenane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X\rangle/L)$, (b) shear viscosity ($\eta$), (c) tumbling frequency ($\omega\tau_{R}$), and (d) first normal stress coefficient ($\psi_{1}$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Colors from cold to hot indicate higher number of rings ($N_{\mathrm{rings}}$) with fixed beads per ring ($\mathrm{BPR}=8$). Fitted power law exponents for $\mathrm{Wi}>10^{2}$ for $\eta$ and $\psi_{1}$ are reported in SI Tables E.1 and E.2, respectively.
@@
Figure 8: Poylcatenane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in flow direction $(\langle X\rangle/L)$, (b) shear viscosity ($\eta$), (c) tumbling frequency ($\omega\tau_{R}$), and (d) first normal stress coefficient ($\psi_{1}$). Darker shades of green indicate increasing beads per ring (BPR) in polycatenanes with fixed number of rings ($N_{\mathrm{rings}}=10)$. Fitted power law exponents for $\mathrm{Wi}>10^{2}$ for $\eta$ and $\psi_{1}$ are reported in SI Tables E.1 and E.2, respectively.